OpenStudy (moongazer):

the ave. of X numbers is 5. If Y is added to each number, the ave. will be?

OpenStudy (anonymous):

Hi moongazer, Average is sum/number, if you increase sum by 5y and divide by 5, then average is 5+y.

OpenStudy (anonymous):

\[\sum_{i}^{X}/X =5\] adding y \[\sum_{i}^{X}/X +y =5+y\] hence average = 5 +y

OpenStudy (moongazer):

waT?

OpenStudy (moongazer):

where did you get 5y?

OpenStudy (anonymous):

5 numbers, 5 y's.

OpenStudy (moongazer):

you are correct but i can't understand how this works

OpenStudy (anonymous):

You added a y to each number and there are 5 numbers so 5y.

OpenStudy (moongazer):

could you give example?

OpenStudy (anonymous):

Example of what?

OpenStudy (anonymous):

1+2+3 = 6, 1+y +2 +y +3 +y = 6 +3y

OpenStudy (moongazer):

ohh ok thanks

OpenStudy (moongazer):

how did you get 5+y?